OpenStudy (anonymous):

solve for the variable in the following equation x/2 - 2/3 =1/4

OpenStudy (anonymous):

x/2 = 1/4+2/3 find LCD x/2 = 3/12 + 8/12 add the right side x/2 = 11/12 multiply by two to get "X" by itself x=(11/12)(2) multiply for your answer x=22/12 now do any reducing you may need to x=11/6 hope it helped